OpenStudy (anonymous):

given r(x)=7x-22 and t(x)=x^2+3, express the sum (r+t)(x).

OpenStudy (akshay_budhkar):

its very easy.. (r+t)(x)=r(x)+t(x) just ad the two functions.

myininaya (myininaya):

just add r and t

OpenStudy (anonymous):

(r+t)(x)=x^2+7x+25

myininaya (myininaya):

-22+3=-19

OpenStudy (akshay_budhkar):

no.. add properly

OpenStudy (akshay_budhkar):

oh kk myininaya did it

myininaya (myininaya):

lol sorry akshay she/he almost had it

OpenStudy (akshay_budhkar):

ya..lol